OpenStudy (anonymous):

(1 / n+ 9) * (3n - 18)/ ( 6 - n)

OpenStudy (anonymous):

\[\frac{ 1 }{ n+9 } * \frac{ 3n-18 }{ 6-n }\]

OpenStudy (anonymous):

factor out the 3 in the numerator in the second fraction

OpenStudy (anonymous):

to get 3 * (n + 9) as the new numerator of the second fraction

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

ask me a question if that first step did not make sense

OpenStudy (anonymous):

you get 3(1-6)

OpenStudy (anonymous):

3(1-n)

OpenStudy (anonymous):

not quite, how did you make the horizontal division lines in your first post? I would like to be able to do that to better help explain

OpenStudy (anonymous):

i clicked the equation blue button

OpenStudy (anonymous):

great, thanks

OpenStudy (anonymous):

okay, factoring out 3 from the numerator (the top part) of the second fraction should look like this:

OpenStudy (anonymous):

\[\frac{ 3 }{ n+9}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ (n + 9) } * \frac{ 3 * (n + 6) }{ (6 - n) }\]

OpenStudy (anonymous):

oops, i made a mistake!

OpenStudy (anonymous):

do you get \[\frac{ 3 }{ n+9 } \] when you done

OpenStudy (anonymous):

close

OpenStudy (anonymous):

\[\frac{ 1 }{ (n+9) } * \frac{ 3 *(n - 6) }{ (6 - n) }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ (n+9) } * \frac{ 3 * (n - 6) }{ - (n - 6) }\]

OpenStudy (anonymous):

\frac{ 1 }{ (n + 9) } * \frac{ 3 }{ -1 }

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 1 }{ (n + 9) } * \frac{ 3 }{ -1 }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ -(n+9) }\]

OpenStudy (anonymous):

is the final answer

OpenStudy (anonymous):

ask me a question if something did not make sense

OpenStudy (anonymous):

how you got a negative on the denometor

OpenStudy (anonymous):

thanks